# Direct Bias-Correction Term Estimation for Propensity Scores and Average Treatment Effect Estimation

### v1

@@ 3.3 Neural networks
This section provides an estimation error analysis when we use neural networks for $\mathcal{R}$}. Our analysis is mostly based on Kato & Teshima (2021) and Zheng et al. (2022).
Feedforward neural networks (FNNs). We define FNNs as follows:
Definition 3.6 (Feedforward neural networks. From Zheng et al. (2022)).
Let $\mathcal{D}$}, $\mathcal{W}$}, $\mathcal{U}$}, and ${\mathcal{S}}\in(0,\infty)$ be parameters that can depend on $n$. Let ${\mathcal{F}}^{\mathrm{FNN}}\coloneqq{\mathcal{F}}^{\mathrm{FNN}}_{M,{\mathcal{D}},{\mathcal{W}},{\mathcal{U}},{\mathcal{S}}}$ be a class of ReLU-activated FNNs $r_{\theta}(1,x)=1/e_{\theta}(x)$ and $r_{\theta}(0,x)=1/(1-e_{\theta}(x))$, where $e_{\theta}\colon{\mathbb{R}}^{K}\to{\mathbb{R}}$ with parameter $\theta$, depth $\mathcal{D}$}, width $\mathcal{W}$}, size $\mathcal{S}$}, number of neurons $\mathcal{U}$}, satisfies the following conditions: (i) the number of hidden layers is $\mathcal{D}$}; (ii) the maximum width of the hidden layers is $\mathcal{W}$}; (iii) the number of neurons in $e_{\theta}$ is $\mathcal{U}$}; (iv) the total number of parameters in $e_{\theta}$ is $\mathcal{S}$}.
Let $\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$ be the pseudo-dimension of ${\mathcal{F}}^{\mathrm{FNN}}$. For the definition, see Anthony & Bartlett (1999) and Definition 3 in Zheng et al. (2022).
We model $r_{0}$ by $r(1,X)=\frac{1}{e(x)}$ and $r(0,X)=\frac{1}{1-e(x)}$, where $e(x)=\frac{1}{1+\exp(-f(x))}$ for $f\in{\mathcal{F}}^{\mathrm{FNN}}$. Let us denote such an estimator by $r_{f}$ for $f\in{\mathcal{F}}^{\mathrm{FNN}}$. Let $f_{0}$ denote the true function such that $r_{f_{0}}=1+\exp(-f_{0}(x))$ holds. Then, the estimator is denoted as $\widehat{f}^{\mathrm{FNN}}_{n}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}^{\mathrm{FNN}}}\widehat{{\mathcal{L}}}_{n}(r_{f})$.
Estimation error analysis. For the estimator, we can prove an estimation error bound. Let us make the following assumption.
@@
Then, we prove the following estimation error bound:
@@
Suppose that Assumption 3.7 holds. Also assume $r_{0}\in\Sigma(\beta,M,[0,1]^{d})$ with $\beta=k+a$, where $k\in{\mathbb{N}}^{+}$ and $a\in(0,1]$, and ${\mathcal{F}}^{\mathrm{FNN}}$ has width $\mathcal{W}$} and depth $\mathcal{D}$} such that ${\mathcal{W}}=38\big(\lfloor\beta\rfloor+1\big)^{2}d^{\lfloor\beta\rfloor+1}$ and ${\mathcal{D}}=21\big(\lfloor\beta\rfloor+1\big)^{2}\lceil n^{\frac{d}{2(d+2\beta)}}\log_{2}\left(8n^{\frac{d}{2(d+2\beta)}}\right)\rceil$. Then, for $M\geq 1$ and $n\leq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, it holds that $\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\left\|e_{0}(X)\left(\widehat{r}^{\mathrm{FNN}}_{n}(1,X)-r_{0}(1,X)\right)\right\|^{2}_{P_{0}}+\left\|(1-e_{0}(X))\left(\widehat{r}^{\mathrm{FNN}}_{n}(0,X)-r_{0}(0,X)\right)\right\|^{2}_{P_{0}}=C_{0}\big(\lfloor\beta\rfloor+1\big)^{9}d^{2\lfloor\beta\rfloor+(\beta\land 3)}n^{-\frac{2\beta}{d+2\beta}}\log^{3}n$, where $C_{0}>0$ is a constant independent of $n$.
@@ 4 Example about the AIPW estimator
We analyze the AIPW estimator with an estimated propensity score. Recall that the AIPW estimator is defined as $\widetilde{\tau}^{\mathrm{AIPW}}_{n}=\frac{1}{n}\sum^{n}_{i=1}\Big(h(X_{i},D_{i};\widehat{r}_{n})\left(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i})\right)+\widehat{\mu}_{n}(1,X_{i})-\widehat{\mu}_{n}(0,X_{i})\Big)$, which is also called the DR estimator.
@@
Assumption 4.1 (Donsker condition).
The hypothesis classes $\mathcal{R}$} and $\mathcal{M}$} belong to the Donsker class.
For example, this assumption holds when the bracketing entropy is finite. In contrast, it is violated in high-dimensional regression or series regression settings where the model complexity diverges as $n\to\infty$. For neural networks, the assumption holds if both the number of layers and the width are finite. However, if these quantities grow with the sample size, the assumption is no longer valid.
Even if this assumption does not hold, we can still establish asymptotic normality by employing sample splitting (Klaassen, 1987). There are various ways to implement sample splitting, and one of the most well-known is cross-fitting, used in double machine learning (DML Chernozhukov et al., 2018, ,). In DML, the dataset is split into several folds, and the nuisance parameters are estimated using only a subset of the folds. This ensures that in $h(X_{i},D_{i};\widehat{r}_{n})\left(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i})\right)+\widehat{\mu}_{n}(1,X_{i})-\widehat{\mu}_{n}(0,X_{i})$, the observations $(X_{i},D_{i},Y_{i})$ are not used to construct $\widehat{\mu}_{n}$ and $\widehat{r}_{n}$. For more details, see Chernozhukov et al. (2018).
@@
$\big\|\widehat{r}-r_{0}\big\|_{2}=o_{p}(1)$, $\big\|\widehat{\mu}-\mu_{0}\big\|_{2}=o_{p}(1)$, and $\big\|\widehat{r}-r_{0}\big\|_{2}\big\|\widehat{\mu}-\mu_{0}\big\|_{2}=o_{p}(1/\sqrt{n})$.
@@
Suppose that Assumptions 1.1, and 4.14.2 hold. Then, the AIPW estimator converges in distribution to a normal distribution as $\sqrt{n}\Big(\widetilde{\tau}^{\mathrm{AIPW}}_{n}-\tau_{0}\Big)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,V^{*})$, where $V^{*}$ is the efficiency bound defined as $V^{*}\coloneqq{\mathbb{E}}\left[\frac{\sigma^{2}(1,X)}{e_{0}(X)}+\frac{\sigma^{2}(0,X)}{1-e_{0}(X)}+\big(\tau_{0}(X)-\tau_{0}\big)^{2}\right]$ and $\tau_{0}(X)\coloneqq{\mathbb{E}}[Y(1)-Y(0)\mid X]$.
@@
Here, $V^{*}$ is the efficiency bound given by the variance of the efficient influence function; that is, $V^{*}={\mathbb{E}}\left[\Psi^{*}(X,D,Y)^{2}\right]$ holds (van der Vaart, 1998). Thus, this estimator is efficient.
5 Generalization via the Bregman divergence minimization
We can further generalize our direct bias-correction term estimation from the viewpoint of Bregman divergence minimization and point out the connection to Riesz regression and covariate balancing.
5.1 Riesz regression and covariate balancing
Our resulting objective function is the same as the one used in Riesz regression (Chernozhukov et al., 2022a, 2024). Although the motivation and derivation differ, we point out that Riesz regression can be interpreted as a specific instance of DRE.
Building on this perspective, we generalize the objective function using Bregman divergence, following Sugiyama et al. (2011). Bregman divergence is a measure of discrepancy between two points, defined in terms of a strictly convex function (Bregman, 1967). In the context of DRE, Sugiyama et al. (2011) demonstrates that various existing methods can be formulated as Bregman divergence minimization problems. Inspired by this idea, we also extend the direct estimation method for the bias-correction term within the Bregman divergence minimization framework.
This generalization further allows us to derive the empirical balancing method proposed in Chan et al. (2015), which aims to achieve covariate balance and known as an instance of tailored loss function (Zhao, 2019). These results suggest that Riesz regression and covariate balancing can be unified under the DRE framework.
5.2 Bregman divergence minimization
Let $\mathcal{G}$} be a set of functions $g\colon{\mathbb{R}}\to{\mathbb{R}}$ that is differentiable and strictly convex. Given $d\in\{1,0\}$, we define the Bregman divergence between $r(d,\cdot),r(d,\cdot)\colon{\mathcal{X}}\to(1,\infty)$ as
$\mathrm{br}^{\dagger}_{g}\big(r_{0}(d,x)\mid r(d,x)\big)\coloneqq g(r_{0}(d,x))-g(r(d,x))-\partial g(r(d,x))(r_{0}(d,x)-r(d,x)),$
where $\partial g$ denotes the derivative of $g$. Then, we define the average Bregman divergence as
$\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(g(r_{0}(d,X))-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big].$
We estimate $r_{0}$ by $r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)$. By dropping the term that is irrelevant to learning, we have
$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}_{g}\big(r\big),$
where $\mathrm{BR}_{g}\big(r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\Big(-g(r(d,X))+\partial g(r(d,X))r(d,X)\Big)-\partial g(r(d,X))\Big]$. For the derivation, see Appendix B.
Then, we estimate the bias-correction term $h_{0}$ (or $r_{0}$) by minimizing an empirical Bregman divergence as
$\widehat{r}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{\mathrm{BR}}_{g}\big(r\big)+\lambda J(r),$
where $\widehat{\mathrm{BR}}_{g}(r)\coloneqq\sum_{d\in\{1,0\}}\frac{1}{n}\sum^{n}_{i=1}\left(\mathbbm{1}[D_{i}=d]\Big(-g(r(d,X_{i}))+\partial g(r(d,X_{i}))r(d,X_{i})\Big)-\partial g(r(d,X_{i}))\right).$
5.3 Least squares
Our least squares method for direct bias-correction term estimation can be obtained by using $g(r)=r^{2}$. Under this choice of $g$, we obtain $\mathrm{BR}_{g}\big(r\big)=\sum_{d\in\{1,0\}}{\mathbb{E}}\left[-2r(d,X)+\mathbbm{1}[D=d]r(d,X)^{2}\right]$. This gives an objective that is the same as the one used in Chernozhukov et al. (2022a).
5.4 Constrained maximum likelihood estimation
By using different choices of $g$, we can derive various objective functions for direct bias-correction term estimation. As a specific case of Bregman divergence minimization, we introduce constrained maximum likelihood estimation.
Consider $g^{\mathrm{L}}(r)=r\log r-r$, which is a convex function. By substituting this function into the Bregman divergence, we obtain
$r^{*}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}_{g^{\mathrm{L}}}\big(r\big).$ (1)
where $\mathrm{BR}_{g^{\mathrm{L}}}\big(r\big)\coloneqq{\mathbb{E}}\big[-\log(r(1,X))-\log(r(0,X))+\mathbbm{1}[D_{i}=1]r(1,X_{i})+\mathbbm{1}[D_{i}=0]r(0,X_{i})\big]$.
Then, we estimate the bias-correction term as $\widehat{r}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{\mathrm{BR}}_{g^{\mathrm{L}}}\big(r\big)+\lambda J(r)$, where $\widehat{\mathrm{BR}}_{g^{\mathrm{L}}}(r)=\frac{1}{n}\sum^{n}_{i=1}\big(-\log\big(r(1,X_{i})\big)-\log\big(r(0,X_{i})\big)+\mathbbm{1}[D_{i}=1]r(1,X_{i})+\mathbbm{1}[D_{i}=0]r(0,X_{i})\big)$. This estimation method corresponds to unnormalized Kullback–Leibler (UKL) minimization in DRE (Sugiyama et al., 2011), which generalizes the KL importance estimation procedure (KLIEP).
Note that solving (1) is equal to
5.5 Empirical balancing
Next, we derive empirical balancing as a special case of Bregman divergence minimization (Chan et al., 2015). Empirical balancing is known as a specific form of covariate balancing and can be derived from a tailored loss function (Zhao, 2019).
Let us consider $g^{\mathrm{E}}(r)=(r-1)\log\left(r-1\right)-r$, which is also convex for $r\in(1,\infty)$ and its derivative is $\partial g^{\mathrm{E}}(r)=\log\left(r-1\right)$. By substituting this function, we obtain $\mathrm{BR}_{g^{\mathrm{E}}}\big(r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\Big(\log\left(r-1\right)+r(d,X)\Big)-\log\left(r-1\right)\Big]$. Note that it holds that $\mathrm{BR}_{g^{\mathrm{E}}}\big(r\big)=$
${\mathbb{E}}\Big[-\mathbbm{1}[D=0]\log\left(r(1,X)-1\right)-\mathbbm{1}[D=1]\log\left(r(0,X)-1\right)+\mathbbm{1}[D=1]r(1,X)+\mathbbm{1}[D=0]r(0,X)\Big].$
Then, we estimate the bias-correction term as $\widehat{r}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{\mathrm{BR}}_{g^{\mathrm{E}}}\big(r\big)$, where the empirical Bregman divergence becomes $\widehat{\mathrm{BR}}_{g^{\mathrm{E}}}(r)=\frac{1}{n}\sum^{n}_{i=1}\big(\mathbbm{1}[D_{i}=0]\log\big(r(1,X_{i})-1\big)+\mathbbm{1}[D_{i}=1]\log\big(r(0,X_{i})-1\big)+\mathbbm{1}[D_{i}=1]r(1,X_{i})+\mathbbm{1}[D_{i}=0]r(0,X_{i})\big)$. If we model $r$ as $r(1,x)=1/e(x)$ and r(0,x)=1/(1e(x))r(0,x)=1/(1-e(x)), we have r(1,x)1=1/(r(0,x)1)r(1,x)-1=1/(r(0,x)-1). Therefore, we have $\widehat{\mathrm{BR}}_{g^{\mathrm{E}}}(r)=$
$\frac{1}{n}\sum^{n}_{i=1}\left(\mathbbm{1}[D_{i}=1]\left(-\log\left(\frac{1}{r(1,X_{i})-1}\right)+r(1,X_{i})\right)+\mathbbm{1}[D_{i}=0]\left(-\log\left(\frac{1}{r(0,X_{i})-1}\right)+r(0,X_{i})\right)\right).$
This objective function is equivalent to the one with the tailored loss proposed in Zhao (2019). From this objective, we can also derive empirical balancing (Chan et al., 2015).
6 Discussion and related work
6.1 Related work
The estimation of the bias-correction term or the propensity score has been a core interest in causal inference. The bias-correction term can be interpreted as a gradient in a one-step estimator (van der Vaart, 2002). This idea was refined by van der Laan (2006) as targeted maximum likelihood estimation (TMLE). Chernozhukov et al. (2024) also explore a related topic from the viewpoint of Riesz representers and double machine learning (DML) (Chernozhukov et al., 2018). They propose automatic DML along with its implementation (Chernozhukov et al., 2022a).
This topic has been addressed from various perspectives. One such approach focuses on estimating the propensity score by matching the distributions of the treatment and control groups (Chan et al., 2015; Deville & Särndal, 1992; Graham et al., 2012; Bryan S. Graham & Egel, 2016; Hellerstein & Imbens, 1999). For example, Imai & Ratkovic (2013) introduces the Covariate Balancing Propensity Score (CBPS), which estimates propensity scores by explicitly balancing covariate means. In parallel with, or subsequent to, Imai & Ratkovic (2013), several methods related to covariate balancing have been proposed (Hainmueller, 2012; Zubizarreta, 2015; Athey et al., 2018).
From a methodological perspective, our study is inspired by DRE (Sugiyama et al., 2012). We review the literature of DRE in Appendix A.
@@ 6.1 Related work
Data Dimension DM DBC (LS) DBC (UKL) Logistic CBPS RieszNet DM

### v2

@@ 4.3 Neural networks
Second, we provide an estimation error analysis when we use neural networks for $\mathcal{H}$}. Our analysis is mostly based on Kato & Teshima (2021) and Zheng et al. (2022). We define Feedforward neural networks (FNNs) as follows:
Definition 4.1 (FNNs. From Zheng et al. (2022)).
Let $\mathcal{D}$}, $\mathcal{W}$}, $\mathcal{U}$}, and ${\mathcal{S}}\in(0,\infty)$ be parameters that can depend on $n$. Let ${\mathcal{F}}^{\mathrm{FNN}}\coloneqq{\mathcal{F}}^{\mathrm{FNN}}_{M,{\mathcal{D}},{\mathcal{W}},{\mathcal{U}},{\mathcal{S}}}$ be a class of ReLU-activated FNNs with parameter $\theta$, depth $\mathcal{D}$}, width $\mathcal{W}$}, size $\mathcal{S}$}, number of neurons $\mathcal{U}$}, satisfies the following conditions: (i) the number of hidden layers is $\mathcal{D}$}; (ii) the maximum width of the hidden layers is $\mathcal{W}$}; (iii) the number of neurons in $e_{\theta}$ is $\mathcal{U}$}; (iv) the total number of parameters in $e_{\theta}$ is $\mathcal{S}$}.
For the model ${\mathcal{F}}^{\mathrm{FNN}}$, we define $\widehat{f}^{\mathrm{FNN}}_{n}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}^{\mathrm{FNN}}}\widehat{{\mathcal{L}}}_{n}(\zeta^{-1}\circ f)$. Then, we define an estimator as $\widehat{h}^{\mathrm{FNN}}_{n}\coloneqq\zeta^{-1}\circ\widehat{f}^{\mathrm{FNN}}_{n}$.
For the estimator, we can prove an estimation error bound. Let us make the following assumption.
@@
Let $\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$ be the pseudo-dimension of ${\mathcal{F}}^{\mathrm{FNN}}$. For the definition, see Anthony & Bartlett (1999) and Definition 3 in Zheng et al. (2022). Then, we prove the following estimation error bound:
@@
Suppose that $g$ is $\mu$-strongly convex and there exist constant $C>0$ such that $|g^{\prime\prime}(t)|\leq M\quad\forall t\in{\mathbb{R}}$. Assume also that $\zeta^{-1}(0)$ is finite. Suppose that Assumption 4.2 holds. For $f_{0}$ such that $h_{0}=\zeta^{-1}\circ f_{0}$, also assume $f_{0}\in\Sigma(\beta,M,[0,1]^{d})$ with $\beta=k+a$, where $k\in{\mathbb{N}}^{+}$ and $a\in(0,1]$, and ${\mathcal{F}}^{\mathrm{FNN}}$ has width $\mathcal{W}$} and depth $\mathcal{D}$} such that ${\mathcal{W}}=38\big(\lfloor\beta\rfloor+1\big)^{2}d^{\lfloor\beta\rfloor+1}$ and ${\mathcal{D}}=21\big(\lfloor\beta\rfloor+1\big)^{2}\lceil n^{\frac{d}{2(d+2\beta)}}\log_{2}\left(8n^{\frac{d}{2(d+2\beta)}}\right)\rceil$. Then, for $M\geq 1$ and $n\leq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, it holds that $\left\|\widehat{h}^{\mathrm{FNN}}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}=C_{0}\big(\lfloor\beta\rfloor+1\big)^{9}d^{2\lfloor\beta\rfloor+(\beta\land 3)}n^{-\frac{2\beta}{d+2\beta}}\log^{3}n$, where $C_{0}>0$ is a constant independent of $n$.
@@ 5 Example about the AIPW estimator
We analyze the AIPW estimator with an estimated propensity score. Recall that the AIPW estimator is defined as $\widetilde{\tau}^{\mathrm{AIPW}}_{n}=\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{h}_{n}(D_{i},X_{i})\left(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i})\right)+\widehat{\mu}_{n}(1,X_{i})-\widehat{\mu}_{n}(0,X_{i})\Big)$, which is also called the DR estimator.
@@
Assumption 5.1 (Donsker condition or cross fitting).
Either of the followings holds: (i) the hypothesis classes $\mathcal{H}$} and $\mathcal{M}$} belong to the Donsker class, or (ii) $\widehat{\mu}_{n}$ and $\widehat{h}_{n}$ are estimated via cross fitting.
For example, the Donsker condition holds when the bracketing entropy of $\mathcal{H}$} is finite. In contrast, it is violated in high-dimensional regression or series regression settings where the model complexity diverges as $n\to\infty$. For neural networks, the assumption holds if both the number of layers and the width are finite. However, if these quantities grow with the sample size, the assumption is no longer valid.
Even if the Donsker condition does not hold, we can still establish asymptotic normality by employing sample splitting (Klaassen, 1987). There are various ways to implement sample splitting, and one of the most well-known is cross-fitting, used in double machine learning (DML, Chernozhukov et al., 2018). In DML, the dataset is split into several folds, and the nuisance parameters are estimated using only a subset of the folds. This ensures that in $\widehat{h}_{n}(D_{i},X_{i})\left(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i})\right)+\widehat{\mu}_{n}(1,X_{i})-\widehat{\mu}_{n}(0,X_{i})$, the observations $(X_{i},D_{i},Y_{i})$ are not used to construct $\widehat{\mu}_{n}$ and $\widehat{r}_{n}$. For more details, see Chernozhukov et al. (2018).
@@
$\big\|\widehat{h}-h_{0}\big\|_{2}=o_{p}(1)$, $\big\|\widehat{\mu}-\mu_{0}\big\|_{2}=o_{p}(1)$, and $\big\|\widehat{h}-h_{0}\big\|_{2}\big\|\widehat{\mu}-\mu_{0}\big\|_{2}=o_{p}(1/\sqrt{n})$.
@@
Suppose that Assumptions 1.1, and 5.15.2 hold. Then, the AIPW estimator converges in distribution to a normal distribution as $\sqrt{n}\Big(\widetilde{\tau}^{\mathrm{AIPW}}_{n}-\tau_{0}\Big)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,V^{*})$, where $V^{*}$ is the efficiency bound defined as $V^{*}\coloneqq{\mathbb{E}}\left[\frac{\sigma^{2}(1,X)}{e_{0}(X)}+\frac{\sigma^{2}(0,X)}{1-e_{0}(X)}+\big(\tau_{0}(X)-\tau_{0}\big)^{2}\right]$ and $\tau_{0}(X)\coloneqq{\mathbb{E}}[Y(1)-Y(0)\mid X]$.
@@
Here, $V^{*}$ matches the efficiency bound given as the variance of the efficient influence function (van der Vaart, 1998). Thus, this estimator is efficient.
@@
Data Dimension DM DBC (LS) DBC (KL) MLE CBPS RieszNet DM